find the center and the radius of the circle
x^2+y^2 =100

Answer by tinbar(128)

circles can be given in the form (x-p)^2+(y-q)^2 = r^2 where (p,q) represent the coordinates of the centre of the circle and r is the radius
your question can be written as (x-0)^2+(y-0)^2=10^2. therefore, your centre is (0,0) and the radius is 10 units